OpenStudy (rose4825):

4. If you were going to make a curtain for a window that was two and a half feet wide and three and a half feet high, what size would your piece of fabric need to be? Add a half foot to both the length and the width to account for seams and give your answer in yards.

OpenStudy (bibby):

we want the area. first add 1/2 a foot to both length and width length+\(\dfrac{1}{2}=\dfrac{1}{2}+\dfrac{1}{2}=1ft\) width+\(\dfrac{1}{2}=2\dfrac{1}{2}+\dfrac{1}{2}==3ft\) We then convert to yards. 1ft = 1/3 yd 3ft = 1 yd, area = len*width=3yd*1/3yd==\(1yd^2\) I'm not positive about this

OpenStudy (rose4825):

thank you